7:30 – 3:30
- Was able to upload a new roster.xml on site
- Got file uploading working here. Need to bring in and test
- Windy day. Still managed a 20.4 at 265 avg watts.
- Meeting – nope, canceled
- Officially requested some monitors
- Added a “custom query” field to the SOLR Manager. Tomorrow I’ll play with accessing the db.
- Need to add pieces to show billet info
- Dong is setting up an HTTPS server. We’ll see if all this stuff still works.